OpenStudy (buttermequeasy):

|3x + 2| = 14 x = 4 and x = −4 x = 4 and x = negative 16 over 3 x = 28 over 3 and x = negative 16 over 3

OpenStudy (anonymous):

3x + 2 = 14, 3x + 2 = -14 3x = 12, 3x = -16 x = 4, -16/3